Determine whether the function has an inverse function and if so, find the inverse function f(x)=x^3+8
Answer by stanbon(75887)

Determine whether the function has an inverse function and if so, find the inverse function f(x)=x^3+8
----------------------------
Interchange x and y to get:
x = y^3+8
Solve for "y":
y^3 = x-8
y = (x-8)^(1/3)
----------------------
Yes, the inverse exists and it is a function.
=============================
Cheers,
Stan H.